# Direct observation of the crystal electric-field splitting under magnetic field and uncovering field-induced magnetic phase transition in triangular rare-earth magnet CsErSe2

### v1

@@ III.2 Magnetization and AC magnetic susceptibility
As shown in Table 1, the exchange energies in CsErSe2   are at the order $10^{-3}$ meV or less, much smaller than both the CEF interaction and the Zeeman splitting. Consequently, the observed magnetic properties, including conspicuous magnetic anisotropy, are governed by the single-ion ground state properties.
@@
Fig. 3 (a) schematically illustrates the two lowest energy levels under field in $\mathbf{B}\parallel c$, where two field scales of the level-crossing are identified: $B_{C}\approx 5.4$ T and $B_{C}^{*}\approx 1.4$ T, where the ground and the first excited state undergo level-crossing.
In order to examine the corresponding macroscopic magnetic properties, we show the measured AC magnetic susceptibility, $\chi^{\rm ac}_{c}$ (solid lines) in Fig. 3 (b) at various $T$’s between 25 mK and 1 K, with offsets. The calculated $\frac{dM_{c}}{dB}$s are shown in dotted lines, normalized by the $B=0$ values at various $T$’s between 25 mK and 1.008K.
@@
Before we discuss the field dependence of $\chi^{\rm ac}_{c}$, we first note the $\chi^{\rm ac}_{c}(T)$ at ZF in Fig. 3(c), which shows a sharp peak at $T_{C}\approx 110$ mK (red solid line). This is the evidence of long-range magnetic order that is not captured by our single-ion picture. To probe this order, we performed elastic neutron scattering at 50 mK, as discussed in Sec.III.3.
@@
Figure 3: (a) Schematic three lowest energy levels under magnetic field in $\mathbf{B}\parallel c$. Two fields scales at which the level crossing occurs at the 1st excited and the ground states are marked as $B_{C}^{*}$ and $B_{C}$, respectively. (b) Measured AC magnetic susceptibility, $\chi^{\rm ac}_{c}$ (solid lines) are plotted as a function of field in $\mathbf{B}\parallel c$ at various $T$’s as indicated. Dotted lines display $\frac{dM_{c}}{dB}$’s at the corresponding temperature using the calculated $M_{c}$ as shown in Fig. 1(e). Both data and calculations at each $T$ are plotted with a constant offset for clarity. Green and pink dashed lines are guides to eyes for the evolution of the features near $B_{C}^{*}$ and $B_{C}$, respectively. (c) $\chi^{\rm ac}_{c}(T)$ measured at 0.05 T as a function of $T$ is shown in the red solid line, where the spontaneous ordering is identified as a peak at $T=110$ mK. The dashed line shows the $T$ dependence of $\frac{dM_{c}}{dB}$ in $B\rightarrow 0$}, which clearly does not capture the long-range ordering. (d) Calculated $M_{c}$ (dotted lines) and the numerical integration of the $\chi^{\rm ac}_{c}(B)$ (solid line) at $T=25$ mK and 830 mK are plotted as shown. The sharp rise of $M_{c}$ at $B=B_{c}$ is significantly muted from 25 mK and 830 mK in the calculation but the estimated $M_{c}$ from $\chi^{\rm ac}_{c}$ exhibits much less changes between these two $T$’s, where the rise of the $M_{c}$ is significantly stifled, compared to the calculation. They are all normalized by the values at $B=12$ T for comparison. Inset: Experimentally measured magnetization $M_{c}$’s (symbols) are plotted together with the calculated one at 2 (blue), 6 (green), and 10 K (red) in solid lines, which agree well.
Now going back to Fig. 3(b), we find the most prominent feature is a strong local maximum at $B_{C}$ in all observed $T$’s. We identify this as the field-induced transition from one eigenstate to another due to the level crossing. The peak height decreases with increasing $T$, becoming barely discernible at 1.008 K. This level crossing is a property of the ground state, an innate characteristic of the single ion Hamiltonian [Eq. (1)], and hence the peak location is independent of $T$ as marked with a vertical pink potted line.
They are also consistent with a strong peak in the derivatives of the calculated $M_{c}$ ($\frac{dM_{c}}{dB}$), although the feature in the data are considerably muted and broadened compared to $\frac{dM_{c}}{dB}$ at all temperatures. This is even more evident in Fig.3 (d), where the integrated $\chi^{\rm ac}_{c}(B)$ at $T=25$ and 830 mK are plotted in comparison to the calculated $M_{c}$. We believe this is attributed to disorders and defects such that not all Er3+are sitting in the identical $D_{3d}$ octahedral environment, resulting in broadening of the CEF levels as well as their field-dependencies. By $T=$ 2 K, the typical value for the lowest temperature available common magnetometer [inset of Fig.3 (d)], the rapid increase of $M_{c}$ near $B_{C}$ completely vanished.
@@
Another prominent feature in $\chi^{\rm ac}_{c}(B)$ is the rapid decrease of $\chi^{\rm ac}_{c}$ in the vicinity of $B_{C}^{*}$. This behavior smoothly moves to the higher field as $T$ increases, which is in contrast with the unwavering peak location centered at $B_{C}$ in Fig. 3 (b). This behavior is explained as a two-level system with non-monotonic $B$-dependent energy gap, $\Delta_{10}(B)$, which has a local maximum $\Delta_{10}(B_{C}^{*})=0.47$ meV as shown in Fig. 3(a). At a low temperature, this is contrasting to the normal $B$-linear Zeeman splitting in paramagnetic state, where the corresponding magnetization monotonic linear increases, $M=\alpha B$, such that the increase of the magnetization with field exhibits two different slopes in $B<B_{C}^{*}$ and $B>B_{C}^{*}$. This feature is clearly displayed in the calculated magnetization shown in Fig. 3(c) with two distinct slopes in $B<B_{C}$. to decrease as $\Delta_{10}$ increases with field, without In fact, when the system is thermalized at a temperature $T\ll\delta(B_{c}^{*})$, the $\chi^{\rm ac}_{c}$ will approach zero as $B\rightarrow B_{c}^{*}$. As increasing $T$, the thermal populations that undertake this crossing at $B_{C}^{*}$ increase exponentially, and this feature becomes broad and the location of $B_{C}^{*}$ moves higher $B$, shown in the green dashed line.
@@
We notice that there is a sharp dip in $\chi^{\rm ac}_{c}(B)$ in the narrow region of $0<B<B_{C}^{*}$, which was not at all captured by the single-ion Hamiltonian. At this point, we do not have any idea what causes this dip at around $B_{d}=0.2$ T with little $T$-dependence, but it is suspected to be linked with the long-range order, based on its rapid disappearance above $T_{C}$ at 110 mK. Within the ordered phase, the sudden change in magnetic susceptibility may indicate a spin-flop transition Blundell (2001). In the next section, we will further examine the magnetic ordering in CsErSe2   with elastic neutron scattering.
@@ III.3 Magnetic structure analysis with neutron scattering
So far, we have shown the anisotropic field evolution of the CEF levels, which enable us to determine the single-ion Hamiltonian with remarkable precision and to make a correct prediction on magnetic behavior. The $g$-factor values are found $g_{\perp}=6.55$ in $\mathbf{B}\parallel b$ and $g_{z}=6.98$, directly calculated from $g_{z(\perp)}=2g_{J}|\langle 0_{\pm}|\hat{J}_{z(\perp)}|0_{\pm(\mp)}\rangle|.$, where $g_{J}$ is Landé $g$-factor of Er3+, $\frac{6}{5}$ and $|0_{\pm}\rangle$ are two doubly degenerate ground states of Eq.(1) $B\rightarrow 0$ limit. Despite the similar $g$-factor values, the magnetic properties are found highly field-orientation-dependent, including the step-like field dependence of the magnetization as a function of field only in $\mathbf{B}\parallel c$ but not in $\mathbf{B}\parallel b$.
@@
Our result reminds us that Zeeman splitting cannot always be assumed to be linear in the applied magnetic field due to two reasons: one is due to a level crossing of the ground state in Fig.3(a), and the other due to a dominant quadratic (second order) component of the Zeeman effect as in the ground state $\mathbf{B}\parallel b$ [Fig. 1(a)] in CsErSe2 . For comparison, the dominant quadratic Zeeman splitting occurs in $\mathbf{B}\parallel c$ for CsYbSe2, a sister compound only with a different RE ionPocs et al. (2021). These two non-linearities make a substantial impact on the $T$ dependence of magnetic susceptibility particularly when $T$ is lower than the first CEF gap at ZF, which is 0.67 meV for CsErSe2 , while 13 meV for CsYbSe2.
One of the most striking features of the field-dependent spectra shown in Fig.2, is multiple crossings and high concentrations of levels within a narrow energy window, i.e., below $E<120$ cm-1. Although the excitations from the ground states alone cannot explain some of the peak positions, narrowly spaced levels imply the thermal population may be responsible for these. In addition, there is an obvious lack of field-dependent spectral lines by three doublets lying above $E>160$ cm-1.
@@
where $\mathcal{H}^{p}_{\rm Z}$ refers to the Zeeman interaction between the spin of Er3+and magnetic field by incoming photon, $\hbar\omega$ is the photon energy, $p_{\Gamma}$ is the Bolzmann thermal population factor of the final state replacing the density of state. $\Delta E$ refer to the energy difference between the initial and final states, $\Delta E=E_{f}-E_{i}$. At a given applied magnetic field, Eq. (7) is directly proportional to $|\langle f|\hat{J}_{\alpha}i\rangle|^{2}$, where $\hat{J}_{\alpha}$ denotes the spin component parallel to external magnetic field Furrerer and Podlesnyak (2006); Schaack. (2000)
@@
While Raman and far-infrared techniques combined with intense magnetic field have long been used to investigate magnetic couplings and magnonic behavior in $d$-electron systems Xiang et al. (2023); Mou et al. (2024); Lujan et al. (2024); Mai et al. (2025), they prove to be equally valuable for resolving crystal electric field (CEF) excitations in 4f compounds. These methods are particularly effective at revealing magnetoelastic interactions, enhanced by strong spin-orbit coupling, including vibronic bound states formed through coupling between optical phonons and CEF excitations Liu et al. (2018); Čermák et al. (2019). Such interactions originating from the mixing of spin and lattice eigenstates Scheie et al. (2025) give rise to new hybridized quasiparticles with field-dependent spectra and dispersions, which are responsible for, for example, unusual thermal-transport phenomena Pocs et al. (2025).
@@ V Summary
We have directly measured the splittings of the doublets CEF levels of CsErSe2 with FIR and Raman spectroscopy under an intense magnetic field, from which we obtain the six Steven’s coefficients and the exchange energies with high precision. We have shown that the magnetic anisotropies of CsErSe2 are prominently manifested in the field-dependent CEF energy spectra: Especially, the ground state under field undergoes level-crossing in $\mathbf{B}\parallel c$ configuration, but not in $\mathbf{B}\parallel b$. This gives rise to a field-induced phase transition in $\mathbf{B}\parallel c$ configuration, where the magnetization abruptly increases at $B_{C}=5.4$ T, manifested as a peak in $\chi^{\rm ac}_{c}(B)$, as the ground state wave function changes across $B_{C}$. Furthermore, the low-lying crossing in the first excited state added the thermally activated drop in $\chi^{\rm ac}_{c}(B)$ at $B_{C}^{*}=1.4$ T, which fades out quickly as increasing $T$.
We also find that CsErSe2 is found to have a spontaneous ordering at $T_{C}=110$ mK in ZF. However, the observed field-orientation dependent magnetic behavior is dictated by a single-ion physics and the spin sector of the Hamiltonian plays a minor or little role in determining macroscopic magnetic properties, even at as low as 25 mK.
Our results show that field-induced level crossings provide a novel route to realizing additional degeneracies, potentially giving rise to unconventional spin ground states. The presence of narrowly spaced CEF levels and their field evolution must also be considered when interpreting field-dependent properties of rare-earth magnets.

### v2

@@ III.2 Magnetization and AC magnetic susceptibility
As shown in Table 1, the exchange energies in CsErSe2   are in the order $10^{-3}$ meV or less, much smaller than both the CEF energy scale and the Zeeman splitting. Consequently, the observed magnetic properties, including conspicuous magnetic anisotropy, are mostly governed by the single-ion ground state [Eq. (1].
@@
Fig. 3 (a) schematically illustrates the two lowest energy levels under field in $\mathbf{B}\parallel c$, where two field scales related to the level-crossings are identified: $B_{C}\approx 5.2$ T and $B_{C}^{*}\approx 1.4$ T, at which the ground and the first excited state undergo a crossing, respectively. In order to examine the corresponding macroscopic magnetic properties, we show the measured AC magnetic susceptibility, $\chi^{\rm ac}_{c}$ (solid lines) in Fig. 3 (b) at various $T$’s between 25 mK and 1008 K, alongside the calculated $\frac{dM_{c}}{dB}$s (dotted lines) normalized by the $B=0$ values.
@@
Before we discuss the field dependence further, we first note the temperature dependence of $\chi^{\rm ac}_{c}$ at the ZF limit in Fig. 3(c): It exhibits a sharp peak at $T_{N}\approx 110$ mK (red solid line). This is evidence of long-range magnetic order that our single-ion picture does not capture. To probe this order, we performed elastic neutron scattering at 50 mK, as discussed in Sec. III.3.
@@
Figure 3: (a) Schematics of two lowest energy levels under magnetic field in $\mathbf{B}\parallel c$, where two field scales marked as $B_{C}$ and $B_{C}^{*}$ refer the level crossings at the ground (blue) and the 1st excited states (red line) respectively. (b) Measured AC magnetic susceptibility, $\chi^{\rm ac}_{c}$ (solid lines) are plotted as a function of field in $\mathbf{B}\parallel c$ at various $T$’s as indicated. Dotted lines display $\frac{dM_{c}}{dB}$ using the calculated $M_{c}(B)$ as shown in Fig. 1(e) Both $\chi^{\rm ac}_{c}$ data and calculated $\frac{dM_{c}}{dB}$ are plotted with a constant offset for clarity. Green and pink dashed lines are guides to the eyes for the evolution of the features near $B_{C}^{*}$ and $B_{C}$, respectively (See text). (c) $\chi^{\rm ac}_{c}(T)$ measured at 0.05 T as a function of $T$ is shown in the red solid line, where the spontaneous ordering is identified as a peak at 110110 mK. The dashed line shows the calculated $T$ dependence of $\frac{dM_{c}}{dB}$ in $B\rightarrow 0$}, which clearly does not capture the long-range ordering. (d) Calculated $M_{c}$ (dotted lines) and the numerical integration of the $\chi^{\rm ac}_{c}(B)$ data (solid line) at $T=25$ mK and 830 mK are plotted . They are all normalized by the values at $B=12$ T for comparison. Inset: Experimentally measured magnetization $M_{c}$’s (symbols) are plotted together with the calculated one at 2 (blue), 6 (green), and 10 K (red) in solid lines, which agree well.
Now going back to Fig. 3(b), we find the most prominent feature here is a strong local maximum at $B_{C}$ in all observed $T$’s. We identify this as the field-induced transition from one eigenstate to another due to the level crossing in the ground state at $B_{C}$. The peak height decreases with increasing $T$ still remains discernible at 1008 mK. This level crossing is a property of the ground state, an innate characteristic of the single ion Hamiltonian [Eq. (1)], and hence the peak position is independent of $T$ as marked with a vertical pink dotted line. The $\chi^{\rm ac}_{c}(B)$ data are consistent with the calculation, the prominent peak at $B_{C}$ in the derivatives of $M_{c}$ ($\frac{dM_{c}}{dB}$), although the features in the data are considerably broadened compared to $\frac{dM_{c}}{dB}$ at all temperatures. This is even more apparent in Fig.3 (d), where the normalized $M_{c}(B)$ curves obtained by integrating the $\chi^{\rm ac}_{c}(B)$ data at $T=25$ and 830 mK are plotted alongside the calculated $M_{c}$. The sharp rise of $M_{c}$ at $B_{C}$ in the calculation is significantly muted as $T$ increases from 25 mK and 830 mK, however, the $M_{c}$ from integrating $\chi^{\rm ac}_{c}$ exhibits much less change between these two $T$’s, where the rise of the $M_{c}$ is also significantly stifled, compared to the calculation, similar brodening seen in $\chi^{\rm ac}_{c}(B)$ in panel (b). These broadening observed in experimental curves is attributed to disorders and defects, which result from the fact that some Er3+ions do not occupy exactly identical $D_{3d}$ octahedral environment, leading to a distribution of CEF level energies and smearing the field-dependent features in the data. When $T$ increases to 2 K and above [inset of Fig.3 (d)], the characteristic features of $M_{c}$ near $B_{C}^{*}$ and $B_{C}$ completely vanished.
@@
Another prominent feature in $\chi^{\rm ac}_{c}(B)$ is the rapid decrease of $\chi^{\rm ac}_{c}$ in the vicinity of $B_{C}^{*}$. This behavior smoothly moves to the higher field as $T$ increases, which is in contrast with the unwavering peak location centered at $B_{C}$ in Fig. 3 (b). This behavior is explained as a two-level system with non-monotonic $B$-dependent energy gap [Fig. 3(a)] with a local maximum $0.47$ meV at $B_{C}^{*}$, where the first excited state crosses with the higher energy level. As the field increases from zero to $B_{C}^{*}$, the susceptibility follows a typical Curie-like behavior (dashed line in Fig. 3(c)) and field-independent magnetic susceptibility. In $B_{C}^{*}<B<B_{C}$, where the Zeeman gap now decreases with increasing field, the rate of the increase of $M_{c}$ with field becomes much slower, despite it is far from saturation. As increasing $T$, the thermal populations that undertake this crossing at $B_{C}^{*}$ increase exponentially, and this feature becomes broadened into the higher field, indicated in the green dashed line. This broadening is characteristics for the level-crossing not in the ground state, but in the adjacent excited state.
@@
We note that there is a sharp dip in $\chi^{\rm ac}_{c}(B)$ in the narrow region of $0<B<B_{C}^{*}$, which becomes significantly weakened above $T_{N}$. This behavior is notnot captured by the single-ion Hamiltonian. At this point, we do not have any clear idea of what is responsible for this dip, nor for its $T$-dependence. We suspect that it is linked with the spontaneous long-range order, based on its rapid disappearance above $T_{N}$ at 110 mK: for example, the sudden depression in magnetic susceptibility within the ordered phase may indicate a spin-flop transition Blundell (2001). In the next section, we will further examine the magnetic ordering in CsErSe2   with elastic neutron scattering.
@@ III.3 Magnetic structure analysis with neutron scattering
We have demonstrated the anisotropic responses of the CEF levels for two magnetic-field orientations. This enables a precise determination of the single-ion Hamiltonian, from which we can accurately predict the magnetic behavior.
The $g$-factor values are found $g_{\perp}=6.55$ in $\mathbf{B}\parallel b$ and $g_{z}=6.98$ in $\mathbf{B}\parallel c$ calculated from $g_{z(\perp)}=2g_{J}|\langle 0_{\pm}|\hat{J}_{z(\perp)}|0_{\pm(\mp)}\rangle|$, where $g_{J}$ of Er3+is $\frac{6}{5}$, and $|0_{\pm}\rangle$ are two doubly degenerate ground states in the $B\rightarrow 0$ limit. Despite similar $g$-factor values, the magnetic properties are found to be highly anisotropic, including a step-like field dependence of magnetization in $\mathbf{B}\parallel c$ but not in $\mathbf{B}\parallel b$.
@@
Our result reminds us that the energy gap between two lowest CEF levels, which is most relevant for the magnetic properties, can have a complex field dependence with the applied magnetic field due to two reasons: one is due to a level crossing of the ground state in Fig.3(a), and the other due to a dominant quadratic (second order) component of the Zeeman splitting Pocs et al. (2021) as exhibited in the ground state $\mathbf{B}\parallel b$ in CsErSe2 [Fig. 1(a)] . For comparison, the dominant quadratic Zeeman splitting occurs in $\mathbf{B}\parallel c$ for CsYbSe2, a sister compound only with a different RE ionPocs et al. (2021). These two non-linearities make a substantial impact on the $T$ dependence of magnetic susceptibility, particularly for temperature lower than the first CEF gap at ZF, which is 0.67 meV for CsErSe2 , while 13 meV for CsYbSe2.
One of the most striking features of the field-dependent spectra shown in Fig.2, is multiple crossings and high concentrations of levels within a narrow energy window in $E<120$ cm-1. Although the excitations from the ground states alone cannot explain some of the peak positions, narrowly spaced levels imply the thermal population may be responsible for these. In addition, there is an obvious lack of field-dependent spectral lines by three doublets lying above $E>160$ cm-1.
@@
where $\mathcal{H}^{p}_{\rm Z}$ refers to the Zeeman interaction between the spin of Er3+and magnetic field by incoming photon, $\hbar\omega$ is the photon energy, $p_{\Gamma}$ is the Boltzmann thermal population factor of the final state replacing the density of states. $\Delta E$ refer to the energy difference between the initial and final states, $\Delta E=E_{f}-E_{i}$. At a given applied magnetic field, Eq. (7) is directly proportional to $|\langle f|\hat{J}_{\alpha}i\rangle|^{2}$, where $\hat{J}_{\alpha}$ denotes the spin component parallel to external magnetic field Furrerer and Podlesnyak (2006); Schaack. (2000)
@@
While Raman and far-infrared techniques combined with intense magnetic fields have long been used to investigate magnetic couplings and magnonic behavior in $d$-electron systems Xiang et al. (2023); Mou et al. (2024); Lujan et al. (2024); Mai et al. (2025), they prove to be equally valuable for resolving crystal electric field (CEF) excitations in 4f compounds. These methods are particularly effective at revealing magnetoelastic interactions, enhanced by strong spin-orbit coupling, including vibronic bound states formed through coupling between optical phonons and CEF excitations Liu et al. (2018); Čermák et al. (2019). Such interactions originating from the mixing of spin and lattice eigenstates Scheie et al. (2025) give rise to new hybridized quasiparticles with field-dependent spectra and dispersions, which are responsible for, for example, unusual thermal-transport phenomena Pocs et al. (2025).
@@ V Summary
We have directly measured the splittings of the Kramers doublets of CsErSe2 with FIR and Raman spectroscopy under a strong magnetic field, from which we obtain the six Stevens’ coefficients and the exchange energies with high precision. We have shown that the magnetic anisotropies of CsErSe2 are prominently manifested in the field-dependent CEF energy spectra: Especially, the ground state under field undergoes level-crossing in $\mathbf{B}\parallel c$ orientation, but not in $\mathbf{B}\parallel b$. This gives rise to a field-induced phase transition, where $M_{c}$ abruptly increases at $B_{C}\approx 5.2$ T, manifested as a peak in $\chi^{\rm ac}_{c}(B)$ by switching the ground state wave function across $B_{C}$. Furthermore, the low-lying crossing in the first excited state added the thermally activated drop in $\chi^{\rm ac}_{c}(B)$ at $B_{C}^{*}=1.4$ T, which fades out quickly as increasing $T$.
We also find that CsErSe2 is found to have a spontaneous magnetic ordering at $T_{N}=110$ mK in ZF. However, the observed field-orientation dependent magnetic behavior is dictated by a single-ion physics and the spin sector of the Hamiltonian plays a minor or little role in determining macroscopic magnetic properties, even at as low as 25 mK.
The presence of narrowly spaced CEF levels and their crossings must also be considered when interpreting field-dependent properties of rare-earth magnets. Therefore, it is imperative to understand the single-ion physics with the well-characterized CEF Hamiltonian before uncovering or applying effective spin models. Our results highlight the capacity of single-ion physics to drive rich and unanticipated phenomena in rare-earth magnetic insulators under applied magnetic field.
Acknowledgements.
We are grateful for the fruitful discussion with Ovi Garlea and Mike Hermele. Experimental work at the University of Colorado Boulder was supported by Award No. DE-SC0021377 of the U.S. Department of Energy (DOE), Basic Energy Sciences, Materials Sciences and Engineering Division (MSE). D.S. acknowledges support from the U.S. Department of Energy (Grant No. DE-FG02-07ER46451) for magneto-Raman measurements. Work at Oak Ridge National Laboratory was supported by the U.S. DOE, MSE. A. S. is supported by the Quantum Science Center (QSC), a National Quantum Information Science Research Center of the U.S. DOE. M.O.A. acknowledges funding from the Laboratory Directed Research & Development Program. The neutron scattering experiments used resources at the High Flux Isotope Reactor, a DOE Office of Science User Facility operated by the Oak Ridge National Laboratory. The beam time was allocated to HB2A on proposal number IPTS-31517.1. A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by the National Science Foundation Cooperative Agreement No. DMR-1644779 and No. DMR-2128556 and the State of Florida.